A rectangle has a length that is twice the width. If the area of rectangle is 98 inches. What would be the perimeter of that piece of board in inches.
Answer by mananth(16946)

let width be x
the length will be 2x
Area = L * W
x*2x = 98
2x^2=98
x^2=49
x=7inches which is the width
7*2 =14 inches will be the lebgth